If the point P given bellow is on the graph of a function f, find the corresponding point on the graph of g.
P=(5,-1),g(x)=f(-3x+2)

Answer by ikleyn(53765)

.

You are given


    f(5) = -1.


It is the exact meaning of this fragment  from the post  "the point P = (5,-1) is on the graph of a function f".


Take  -3x+2 = 5,  i.e.  3x = 2-5 = -3,   or  x = -1.

Then  

      g(-1) = f(-3x+2) = f(3+2) = f(5) = (at this point recall that f(5) = -1, so you can continue the equality) = -1.


Thus the corresponding point on the graph of the function g  is  Q = (-1,-1).


ANSWER.  The corresponding point on the graph of the function g  is  Q = (-1,-1).

------------

The lesson to learn from my post is THIS :

    The "corresponding point" in this context is the point, whose coordinates you can calculate, 
    starting from the given point P and using the given definition of the function g.